# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 3.2.3 Existence and uniqueness
The proof consists of several steps.
The vanishing viscosity limit as $\gamma\to 0^{+}$. Thanks to the uniform estimates (3.42), (3.49), (3.63), (3.64), (3.65), (3.70), (3.71) and (3.72) obtained in the previous section, we are able to first extend the unique strong solution $(\varphi^{\gamma,k},\mu^{\gamma,k},\sigma^{\gamma,k})$ of the approximate problem (3.36a)–(3.38) from $[0,T]$ to the whole interval $[0,+\infty)$. Then we can pass to the limit as $\gamma\to 0^{+}$ such that there exists a convergent subsequence $\{(\varphi^{\gamma,k},\mu^{\gamma,k},\sigma^{\gamma,k})\}$ (not relabelled for simplicity) with the limit denoted by $(\varphi^{k},\mu^{k},\sigma^{k})$. The procedure follows a standard compactness argument (see e.g., [2, 21]) and its details are omitted here. The limit triple $(\varphi^{k},\mu^{k},\sigma^{k})$ satisfies
$\varphi^{k}\in L^{\infty}(0,T;W^{2,6}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega)),$
$\mu^{k}\in L^{\infty}(0,T;H^{1}(\Omega)),$
$\sigma^{k}\in L^{\infty}(0,T;H^{1}(\Omega))\cap L^{2}(0,T;H^{2}(\Omega))\cap H^{1}(0,T;L^{2}(\Omega)),$
$\varphi^{k}\in L^{\infty}(\Omega\times(0,T))\ \ \textrm{with}\ \ |\varphi^{k}|<1\quad\textrm{a.e.\ in}\ \Omega\times(0,T),$
for any $T>0$. It solves the system
$\partial_{t}\varphi^{k}+\bm{v}^{k}\cdot\nabla\varphi^{k}=\Delta\mu^{k}-\alpha(\overline{\varphi^{k}}-c_{0}),$ (3.73)
$\mu^{k}=-\Delta\varphi^{k}+\varPsi^{\prime}(\varphi^{k})-\chi\sigma^{k}+\beta\mathcal{N}(\varphi^{k}-\overline{\varphi^{k}}),$
$\partial_{t}\sigma^{k}+\bm{v}^{k}\cdot\nabla\sigma^{k}=\Delta(\sigma^{k}-\chi\varphi^{k}),$
almost everywhere in $\Omega\times(0,+\infty)$, and fulfills the boundary and initial conditions
${\partial}_{\bm{n}}\varphi^{k}={\partial}_{\bm{n}}\mu^{k}={\partial}_{\bm{n}}\sigma^{k}=0,\qquad\qquad$ a.e. on $\ \partial\Omega\times(0,+\infty),$
$\varphi^{k}|_{t=0}=\varphi_{0,k}(x),\ \ \sigma^{k}|_{t=0}=\sigma_{0,k}(x),\qquad$ a.e. in $\ \Omega.$
Passage to the limit as $k\to+\infty$. For the solution $(\varphi^{k},\mu^{k},\sigma^{k})$ obtained in the previous step, we can repeat the argument in Section 3.2.2 without the viscous term (i.e., $\gamma=0$), and show that estimates similar to (3.39), (3.40), (3.42), (3.49), (3.58), (3.61) are still valid, being uniformly with respective to the parameter $k$. In particular, the energy identity holds
$\mathcal{F}^{k}(T)+\int_{0}^{T}\mathcal{D}^{k}(s)\,\mathrm{d}s$
$\quad=\mathcal{F}^{k}(0)-\alpha\int_{0}^{T}\big{(}\overline{\varphi^{k}}-c_{0}\big{)}\int_{\Omega}\mu^{k}\,\mathrm{d}x\,\mathrm{d}s-\int_{0}^{T}\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{k})\mu^{k}\,\mathrm{d}x\,\mathrm{d}s$
$\qquad+\chi\int_{0}^{T}\int_{\Omega}(\bm{v}^{k}\cdot\nabla\sigma^{k})\varphi^{k}\,\mathrm{d}x\,\mathrm{d}s=:\mathcal{R}^{k},$
where $\mathcal{F}^{k}$ and $\mathcal{D}^{k}$ are defined as in (3.17), (3.18), respectively. On the other hand, integrating (3.41) on $[0,T]$, we find
$\mathcal{F}^{\gamma,k}(T)+\int_{0}^{T}\widetilde{\mathcal{D}}^{\gamma,k}(s)\,\mathrm{d}s$
$\quad=\mathcal{F}^{\gamma,k}(0)-\alpha\int_{0}^{T}\big{(}\overline{\varphi^{\gamma,k}}-c_{0}\big{)}\int_{\Omega}\mu^{\gamma,k}\,\mathrm{d}x\,\mathrm{d}s-\int_{0}^{T}\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\mu^{\gamma,k}\,\mathrm{d}x\,\mathrm{d}s$
$\qquad+\chi\int_{0}^{T}\int_{\Omega}(\bm{v}^{k}\cdot\nabla\sigma^{\gamma,k})\varphi^{\gamma,k}\,\mathrm{d}x\,\mathrm{d}s=:\mathcal{R}^{\gamma,k}.$
Here, we focus on the convergent subsequence $\big{\{}(\varphi^{\gamma,k},\mu^{\gamma,k},\sigma^{\gamma,k})\big{\}}$ (not relabelled for simplicity) with the limit $(\varphi^{k},\mu^{k},\sigma^{k})$ and emphasize that all the convergence results obtained below should be understood in the sense of a subsequence. Clearly, we have $\mathcal{F}^{\gamma,k}(0)=\mathcal{F}^{k}(0)$ and it is straightforward to check that as $\gamma\to 0^{+}$,
$\mathcal{F}^{\gamma,k}(T)\to\mathcal{F}^{k}(T),\qquad\mathcal{R}^{\gamma,k}\to\mathcal{R}^{k},$
$\int_{0}^{T}\|\nabla(\sigma^{\gamma,k}(s)-\chi\varphi^{\gamma,k}(s))\|^{2}\,\mathrm{d}s\to\int_{0}^{T}\|\nabla(\sigma^{k}(s)-\chi\varphi^{k}(s))\|^{2}\,\mathrm{d}s.$
Like in [2, Section 2], we can take the difference of the above two energy identities and conclude that
$\lim_{\gamma\to 0^{+}}\int_{0}^{T}\big{(}\|\nabla\mu^{\gamma,k}(s)\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(s)\|^{2}\big{)}\,\mathrm{d}s=\int_{0}^{T}\|\nabla\mu^{k}(s)\|^{2}\,\mathrm{d}s.$
Hence, by the lower semi-continuity of norms with respect to the weak convergence, we can pass to the limit as $\gamma\to 0^{+}$ in the inequalities (3.70), (3.71) to get (with the aid of (3.72))
$\|\nabla\mu^{k}(t)\|^{2}$ $\leq 4\left(\|\nabla\mu_{0,k}\|^{2}+C\alpha|\overline{\varphi_{0,k}}-c_{0}|+\int_{0}^{t}\mathcal{H}^{k}_{2}(s)\,\mathrm{d}s\right)$
×$\quad\times\exp\left(C\int_{0}^{t}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s\right)+4C_{6}(\alpha+\alpha^{2})|\overline{\varphi^{k}(t)}-c_{0}|,\quad\forall\,t\in[0,T],$ (3.74)
and
$\int_{0}^{T}\|\nabla\partial_{t}\varphi^{k}(s)\|^{2}\,\mathrm{d}s$
$\quad\leq 2\left(\|\nabla\mu_{0,k}\|^{2}+C\alpha|\overline{\varphi_{0,k}}-c_{0}|\right)+\left(\sup_{t\in[0,T]}\|\nabla\mu^{k}(t)\|^{2}+4C_{6}(\alpha+\alpha^{2})\right)\int_{0}^{T}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s$
$\qquad+2\int_{0}^{T}\mathcal{H}^{k}_{2}(s)\,\mathrm{d}s+4C_{6}(\alpha+\alpha^{2})|\overline{\varphi^{k}(T)}-c_{0}|,$ (3.75)
where
$\mathcal{H}^{k}_{2}(t)$ $=C(\alpha+\alpha^{2})|\overline{\varphi^{k}(t)}-c_{0}|\|\nabla\bm{v}^{k}(t)\|^{2}+C(\alpha^{2}+\alpha^{4})|\overline{\varphi^{k}(t)}-c_{0}|$
$\quad+C\big{(}\|\nabla\bm{v}^{k}(t)\|^{2}+\|\nabla\mu^{k}(t)\|^{2}+\|\nabla(\sigma^{k}(t)-\chi\varphi^{k}(t))\|^{2}\big{)}.$ (3.76)
In summary, we find the following estimates
$\|\varphi^{k}\|_{L^{\infty}(0,+\infty;W^{2,6}(\Omega))}+\|\partial_{t}\varphi^{k}\|_{L^{2}(0,+\infty;H^{1}(\Omega))}+\|\mu^{k}\|_{L^{\infty}(0,+\infty;H^{1}(\Omega))}$
$\quad+\|\sigma^{k}\|_{L^{\infty}(0,+\infty;L^{6}(\Omega))}+\|\partial_{t}\sigma^{k}\|_{L^{2}(0,+\infty;H^{1}(\Omega)^{\prime})}+\|\nabla(\sigma^{k}-\chi\varphi^{k})\|_{L^{2}(0,+\infty;\bm{L}^{2}(\Omega))}$
$\quad+\|\varPsi^{\prime}(\varphi^{k})\|_{L^{\infty}(0,+\infty;L^{6}(\Omega))}\leq C_{7}.$ (3.77)
Moreover, since
$\|\bm{v}^{k}\cdot\nabla\varphi^{k}\|_{L^{2}(0,T;H^{1}(\Omega))}$
$\quad\leq\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{H}^{1}(\Omega))}\|\nabla\varphi^{k}\|_{L^{\infty}(0,T;L^{\infty}(\Omega))}+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{6}(\Omega))}\|\varphi^{k}\|_{L^{\infty}(0,T;W^{2,3}(\Omega))},$
we easily obtain by the elliptic regularity theory applied to (3.73) that
$\|\mu^{k}\|_{L^{2}(0,T;H^{3}(\Omega))}\leq C_{8}(1+T).$
Here, the positive constants $C_{7},C_{8}$ depend on $\|\widetilde{\mu}_{0}\|_{H^{1}}$, $\|\varphi_{0}\|_{H^{1}}$, $\|\sigma_{0}\|_{H^{1}}$, $\max_{r\in[-1,1]}|\varPsi(r)|$, coefficients of the system, $\Omega$ and in particular, $\|\bm{v}\|_{L^{2}(0,T;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))}$, but they are independent of $k$ and $T$.
With the above uniform estimates, we can apply the standard compactness argument and pass to the limit as $k\to+\infty$. There exists a convergent subsequence $\big{\{}(\varphi^{k},\mu^{k},\sigma^{k})\}$ whose limit is denoted by $(\varphi,\mu,\sigma)$. It is straightforward to check that the triple $(\varphi,\mu,\sigma)$ is a solution to the original problem (3.1)–(3.3) on $[0,+\infty)$ with the corresponding regularity properties as stated in Theorem 3.2.
@@
Uniqueness and further properties. Uniqueness of solution follows from the fact $\sigma\in L^{\infty}(0,T;L^{6}(\Omega))\subset L^{8}(0,T;L^{4}(\Omega))$ for any $T>0$ and the conditional uniqueness result presented in Theorem 3.1-(2). Moreover, the energy identity (3.8) can be obtained by testing (3.1a) with $\mu$, (3.1c) with $\sigma-\chi\varphi$, adding the resultants together and integrating over $[t_{1},t_{2}]$. By comparison in (3.1a), we find
Thus, if we assume in addition, $\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$, then it holds $\partial_{t}\varphi\in L^{\infty}(0,+\infty;(H^{1}(\Omega))^{\prime})$.
@@
Passing to the limit as $k\to+\infty$ in (3.74)–(3.76) and then letting $T\to+\infty$, we have the following explicit bounds:
Corollary 3.1.
Let $(\varphi,\mu,\sigma)$ be a global solution to problem (3.1)–(3.3) obtained in Theorem 3.2. Then it holds
$\sup_{t\in[0,+\infty)}\|\nabla\mu(t)\|^{2}$ $\leq 4\left(\|\nabla\widehat{\mu}_{0}\|^{2}+C\alpha|\overline{\varphi_{0}}-c_{0}|+\int_{0}^{\infty}\mathcal{H}(s)\,\mathrm{d}s\right)$
×$\quad\times\exp\left(C\int_{0}^{\infty}\|\nabla\bm{v}(s)\|^{2}\,\mathrm{d}s\right)+C(\alpha+\alpha^{2}),$
and
$\int_{0}^{\infty}\|\nabla\partial_{t}\varphi(s)\|^{2}\,\mathrm{d}s$ $\leq 2\left(\|\nabla\widehat{\mu}_{0}\|^{2}+C\alpha|\overline{\varphi_{0}}-c_{0}|\right)$
$\quad+\left(\sup_{t\in[0,+\infty)}\|\nabla\mu(t)\|^{2}+C(\alpha+\alpha^{2})\right)\int_{0}^{\infty}\|\nabla\bm{v}(s)\|^{2}\,\mathrm{d}s$
$\quad+2\int_{0}^{\infty}\mathcal{H}(s)\,\mathrm{d}s+C(\alpha+\alpha^{2}),$
where $\widehat{\mu}_{0}=-\Delta\varphi_{0}+\varPsi^{\prime}(\varphi_{0})-\chi\sigma_{0}+\beta\mathcal{N}(\varphi_{0}-\overline{\varphi_{0}})$ and
$\mathcal{H}(t)$ $=C(\alpha+\alpha^{2})|\overline{\varphi(t)}-c_{0}|\|\nabla\bm{v}(t)\|^{2}+C(\alpha^{2}+\alpha^{4})|\overline{\varphi(t)}-c_{0}|$
$\quad+C\big{(}\|\nabla\bm{v}(t)\|^{2}+\|\nabla\mu(t)\|^{2}+\|\nabla(\sigma(t)-\chi\varphi(t))\|^{2}\big{)}.$ (3.78)
The positive constant $C$ depends on $\|\varphi_{0}\|_{H^{1}}$, $\|\sigma_{0}\|_{H^{1}}$, $\overline{\varphi_{0}}$, coefficients of the system and $\Omega$.
@@ 4 Proof of Main Results
$\overline{\varphi}(t)-c_{0}=(\overline{\varphi_{0}}-c_{0})e^{-\alpha t}\quad\text{as well as}\quad\overline{\sigma}(t)=\overline{\sigma_{0}},\quad\forall\,t\geq 0.$ (4.2)
@@
where $C>0$ depends on $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, coefficients of the system and $\Omega$. This together with (4.1), (H1), Korn’s inequality yields the uniform-in-time estimate:
@@
$\quad+\int_{0}^{\infty}\left(\|\nabla\bm{v}(t)\|^{2}+\|\nabla\mu(s)\|^{2}+\|\nabla(\sigma(s)-\chi\varphi(s))\|^{2}\right)\mathrm{d}s\leq C.$ (4.3)
In what follows, we characterize the regularizing effect of the global weak solution at different stages.
@@ 4.1 Instantaneous regularity of $(\varphi,\mu,\sigma)$
We start with the first part of Theorem 2.1.
@@
Thanks to the properties $\mu\in L^{2}(0,1;H^{1}(\Omega))$, $\varphi\in L^{4}(0,1;H^{2}_{N}(\Omega))$ and moreover, $\sigma\in L^{2}(0,1;H^{1}(\Omega))$, then for any $\tau\in(0,1)$, there exists $\widetilde{\tau}\in(0,\tau)$ such that
@@
with $\|\varphi(\widetilde{\tau})\|_{L^{\infty}}\leq 1$ and $\overline{\varphi(\widetilde{\tau})}\in(-1,1)$. Taking $\widetilde{\tau}$ as the initial time, we apply Theorem 3.2 on the interval $[\widetilde{\tau},+\infty)$. This combined with the conditional uniqueness result in Theorem 3.1-(2) allow us to achieve the conclusion of Theorem 2.1-(1). ∎
@@ 4.2 The $\omega$-limit set and eventual separation property of $\varphi$
Like in Section 4.1, there exists $\widetilde{\tau}\in(0,1)$ such that we can apply Theorem 3.2 on $[\widetilde{\tau},+\infty)$ and apply (3.77) to conclude
$\sup_{t\geq 1}\left(\|\varphi(t)\|_{W^{2,6}}+\|\sigma(t)\|_{L^{6}}\right)+\int_{1}^{\infty}\|\partial_{t}\varphi(s)\|_{H^{1}}^{2}\,\mathrm{d}s+\int_{1}^{\infty}\|\partial_{t}\sigma(s)\|_{(H^{1})^{\prime}}^{2}\,\mathrm{d}s\leq C.$ (4.4)
@@
According to (4.2), if $\alpha=0$, we have the mass conservation $\overline{\varphi(t)}=\overline{\varphi_{0}}$ for all $t\geq 0$, if $\alpha>0$, then $\overline{\varphi(t)}$ converges to $c_{0}$ as $t\to+\infty$. In what follows, we focus on the more involved case with $\alpha>0$. The case $\alpha=0$ can be treated in a similar manner with minor modifications.
@@
$\omega(\bm{v},\varphi,\sigma)$ =$=\big{\{}(\bm{v}_{\infty},\varphi_{\infty},\sigma_{\infty})\in\mathcal{Z}:\ \exists\,\left\{t_{n}\right\}\nearrow+\infty\text{ such that }$
@@
Next, we prove (4.5) by adapting the argument for [2, Lemma 3.2]. Using the energy inequality (2.3) for the full system and the energy equality (3.8) for the subsystem of $(\varphi,\sigma)$ (for $t\geq 1$), we can deduce that
$\frac{1}{2}\|\bm{v}(t_{2})\|^{2}+\int_{t_{1}}^{t_{2}}\!\int_{\Omega}2\nu(\varphi)|D\bm{v}|^{2}\mathrm{d}x\mathrm{d}s\leq\frac{1}{2}\|\bm{v}(t_{1})\|^{2}-\int_{t_{1}}^{t_{2}}\!\int_{\Omega}\bm{v}\cdot\nabla(\mu+\chi\sigma)\varphi\,\mathrm{d}x\mathrm{d}s,$
for almost all $t_{1}\geq 1$ and all $t_{2}\in[t_{1},+\infty)$. Using the incompressibility condition and (4.1), the second term on the right-hand side can be controlled as follows:
$-\int_{t_{1}}^{t_{2}}\!\int_{\Omega}\bm{v}\cdot\nabla(\mu+\chi\sigma)\varphi\,\mathrm{d}x\mathrm{d}s\leq\int_{t_{1}}^{t_{2}}\!\int_{\Omega}|\bm{v}\cdot\nabla(\mu+\chi\sigma)\varphi|\,\mathrm{d}x\mathrm{d}s$
$\quad\leq(\|\nabla\mu\|_{L^{2}(t_{1},+\infty;\bm{L}^{2}(\Omega))}+|\chi|\|\nabla(\sigma-\chi\varphi)\|_{L^{2}(t_{1},+\infty;\bm{L}^{2}(\Omega))})\|\bm{v}\|_{L^{2}(t_{1},+\infty;\bm{L}^{2}(\Omega))},$ (4.7)
@@
$\sup_{t\geq t_{1}(\varepsilon)}\frac{1}{2}\|\bm{v}(t)\|^{2}+\int_{t_{1}(\varepsilon)}^{\infty}\!\int_{\Omega}2\nu(\varphi(\tau))|D\bm{v}(s)|^{2}\,\mathrm{d}x\mathrm{d}s\leq\frac{1}{2}\varepsilon^{2}+(1+|\chi|)\varepsilon^{2},$
@@
$\alpha\int_{0}^{\infty}\!\int_{\Omega}|\overline{\varphi}-c_{0}||\mu|\,\mathrm{d}x\mathrm{d}s\leq C\alpha\int_{0}^{\infty}e^{-\alpha s}(1+\|\nabla\mu\|)\,\mathrm{d}s\leq C\alpha\left(1+\int_{0}^{\infty}\|\nabla\mu\|^{2}\,\mathrm{d}s\right)<\infty,$
@@
By interpolation, we infer from (4.4) that $\varphi\in BUC([1,+\infty);W^{2-\epsilon,6}(\Omega))$ for any $\epsilon\in(0,1/2)$. Thanks to the compactness embedding and uniqueness of the weak and strong limit, it holds $\varphi(t_{n})\to\varphi_{\infty}$ strongly in $W^{2-\epsilon,6}(\Omega)$ as $t_{n}\to+\infty$. Recalling the instantaneous regularity of $(\varphi,\mu,\sigma)$, for $t\geq 1$, we are allowed to test (2.1b) by $-\chi(\sigma-\chi\varphi)$ and (2.1c) by $\sigma-\chi\varphi$, respectively. Adding the resultants together yields
@@
$\lim_{t_{n}\to+\infty}\|\sigma(t_{n})-[\chi\varphi_{\infty}+(\overline{\sigma_{0}}-\chi c_{0})\|=0,$
and thus $\sigma_{\infty}=\chi\varphi_{\infty}+(\overline{\sigma_{0}}-\chi c_{0})\in W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega)$.

### v2

@@ 3.2.2 Decoupled auxiliary problem for $(\varphi,\mu)$
Let the assumptions of Theorem 3.2 be satisfied.
Step 1. We can apply Theorem 3.1 to conclude that problem (3.1)–(3.3) with the given velocity field $\bm{v}$ admits a global weak solution $(\varphi^{*},\mu^{*},\sigma^{*})$ on $0,+\infty)$.
Step 2. According to Proposition 3.1-(3), the decoupled auxiliary problem (3.34) with $\varphi=\varphi^{*}$, the same velocity field $\bm{v}$ and the same initial datum $\sigma_{0}\in H^{1}(\Omega)\hookrightarrow L^{6}(\Omega)$ as for problem (3.1)–(3.3), admits a unique global weak solution $\sigma^{\sharp}$ on $0,+\infty)$ such that
$\sigma^{\sharp}\in L^{\infty}(0,+\infty;L^{6}(\Omega))\cap L_{\mathrm{uloc}}^{2}([0,+\infty);H^{1}(\Omega))\cap H^{1}_{\mathrm{uloc}}([0,+\infty);(H^{1}(\Omega))^{\prime}),$
and $\sigma^{\sharp}(0)=\sigma_{0}\in H^{1}(\Omega)$. On the other hand, we easily check that $\sigma^{*}$ is actually a global weak solution to the same problem (3.34) with properties described as in Proposition 3.1-(1). Hence, by the conditional uniqueness result (that is, Proposition 3.1-(2)), we find $\sigma^{*}=\sigma^{\sharp}$ on $0,+\infty)$.
Step 3. Consider now the decoupled auxiliary problem (3.55) with $\sigma=\sigma^{\sharp}$, the same velocity field $\bm{v}$ and the same initial datum $\varphi_{0}$ as for problem (3.1)–(3.3). According to Proposition 3.2-(2), problem (3.55) admits a unique global strong solution $(\varphi^{\sharp},\mu^{\sharp})$ on $0,+\infty)$. On the other hand, we can check that $(\varphi^{*},\mu^{*})$ is indeed a global weak solution to the same problem (3.55) with properties described as in Proposition 3.2-(1). Due to the uniqueness of weak solutions, we have $(\varphi^{*},\mu^{*})=(\varphi^{\sharp},\mu^{\sharp})$ on $0,+\infty)$.
In summary, we have shown $(\varphi^{*},\mu^{*},\sigma^{*})=(\varphi^{\sharp},\mu^{\sharp},\sigma^{\sharp})$ on $0,+\infty)$. In addition, by a comparison in (3.1c), we see that
$\int_{0}^{+\infty}\|\partial_{t}\sigma^{*}(s)\|_{(H^{1})^{\prime}}^{2}\,\mathrm{d}s$ $\leq C\int_{0}^{+\infty}\big(\|\bm{v}(s)\|_{\bm{L}^{6}}^{2}\|\sigma^{*}(s)\|_{L^{3}}^{2}+\|\nabla(\sigma^{*}(s)-\chi\varphi^{*}(s))\|^{2}\big)\,\mathrm{d}s$
$\leq C\sup_{t\geq 0}\|\sigma^{*}(s)\|_{L^{3}}^{2}\int_{0}^{+\infty}\|\bm{v}(s)\|_{\bm{H}^{1}}^{2}\,\mathrm{d}s$
$\quad+C\int_{0}^{+\infty}\|\nabla(\sigma^{*}(s)-\chi\varphi^{*}(s))\|^{2}\,\mathrm{d}s,$
@@
which implies that $\sigma^{*}\in H^{1}([0,+\infty);(H^{1}(\Omega))^{\prime})$. Thanks to the regularity properties of $\sigma^{\sharp}$ and $(\varphi^{\sharp},\mu^{\sharp})$ guaranteed by Proposition 3.1-(3) and Proposition 3.2-(2), respectively, we observe that $(\varphi^{*},\mu^{*},\sigma^{*})$ is actually a global solution of the target problem (3.1)–(3.3) on $0,+\infty)$ with the regularity properties described in Theorem 3.2. Its uniqueness follows from the conditional uniqueness result established in Theorem 3.1-(2). Finally, the regularity of $(\varphi^{*},\mu^{*},\sigma^{*})$ enables us to derive the energy identity (3.7) by testing (3.1a) with $\mu$ and (3.1c) with $\sigma-\chi\varphi$, respectively, adding the resultants together, integrating over $\Omega$ and then $t_{1},t_{2}]$.
@@ 4 Proof of Main Results
$\overline{\varphi}(t)-c_{0}=(\overline{\varphi_{0}}-c_{0})e^{-\alpha t},\quad\overline{\sigma}(t)=\overline{\sigma_{0}},\quad\forall\,t\geq 0.$ (4.2)
@@
where $C>0$ depends on $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, the coefficients of the system, and $\Omega$. This together with (4.1), (H1), Korn’s inequality yields the uniform-in-time estimate:
@@
$\quad+\int_{0}^{+\infty}\left(\|\nabla\bm{v}(t)\|^{2}+\|\nabla\mu(s)\|^{2}+\|\nabla(\sigma(s)-\chi\varphi(s))\|^{2}\right)\mathrm{d}s\leq C.$ (4.3)
In what follows, we characterize the regularizing effect of the global weak solution at different time stages.
@@ 4.1 Instantaneous regularity of $(\varphi,\mu,\sigma)$
We begin by proving the first part of Theorem 2.1.
@@
Due to the regularity properties $\mu\in L^{2}(0,1;H^{1}(\Omega))$, $\varphi\in L^{4}(0,1;H^{2}_{N}(\Omega))$ and $\sigma\in L^{2}(0,1;H^{1}(\Omega))$, for any $\tau\in(0,1)$, there exists $\widetilde{\tau}\in(0,\tau)$ such that
@@
[…]
@@ 4.2 The $\omega$-limit set and eventual separation property of $\varphi$
[…]
$\|\varphi\|_{L^{\infty}(1,+\infty;W^{2,6}(\Omega))}+\|\sigma\|_{L^{\infty}(1,+\infty;L^{6}(\Omega))}\leq C.$ (4.4)
In addition, we have
$\int_{1}^{+\infty}\|\partial_{t}\varphi(s)\|_{(H^{1})^{\prime}}^{2}\,\mathrm{d}s+\int_{1}^{+\infty}\|\partial_{t}\sigma(s)\|_{(H^{1})^{\prime}}^{2}\,\mathrm{d}s\leq C.$ (4.5)
@@
According to (4.2), if $\alpha=0$, we have the mass conservation $\overline{\varphi(t)}=\overline{\varphi_{0}}$ for all $t\geq 0$, if $\alpha>0$, then $\overline{\varphi(t)}$ converges exponentially fast to $c_{0}$ as $t\to+\infty$. In what follows, we focus on the more involved case with $\alpha>0$. The case $\alpha=0$ can be treated in a similar manner with minor modifications.
@@
$\omega(\bm{v},\varphi,\sigma)$ $=\big\{(\bm{v}_{\infty},\varphi_{\infty},\sigma_{\infty})\in\mathcal{Z}:\ \exists\,\left\{t_{n}\right\}\nearrow+\infty\text{ such that }$
@@
Next, we prove (4.6) by adapting the argument for [3, Lemma 3.2]. Using the energy inequality (2.5) for the full system and the energy equality (3.7) for the subsystem of $(\varphi,\sigma)$ (for $t\geq 1$), we can deduce that
$\frac{1}{2}\|\bm{v}(t_{2})\|^{2}+\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}2\nu(\varphi(s))|D\bm{v}(s)|^{2}\,\mathrm{d}x\mathrm{d}s$
$\quad\leq\frac{1}{2}\|\bm{v}(t_{1})\|^{2}-\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}\big[\bm{v}(s)\cdot\nabla(\mu(s)+\chi\sigma(s))\big]\varphi(s)\,\mathrm{d}x\mathrm{d}s,$
for almost all $t_{1}\geq 1$ and all $t_{2}\in[t_{1},+\infty)$. Using the incompressibility condition $\mathrm{div}\bm{v}=0$ and (4.1), the second term on the right-hand side can be controlled as follows:
$-\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}\big[\bm{v}(s)\cdot\nabla(\mu(s)+\chi\sigma(s))\big]\varphi(s)\,\mathrm{d}x\mathrm{d}s$
$\quad=-\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}(\bm{v}(s)\cdot\nabla\mu(s))\varphi(s)\,\mathrm{d}x\mathrm{d}s-\chi\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}\big[\bm{v}(s)\cdot\nabla(\sigma(s)-\chi\varphi(s))\big]\varphi(s)\,\mathrm{d}x\mathrm{d}s$
$\quad\leq\big(\|\nabla\mu\|_{L^{2}(t_{1},+\infty;\bm{L}^{2}(\Omega))}+|\chi|\|\nabla(\sigma-\chi\varphi)\|_{L^{2}(t_{1},+\infty;\bm{L}^{2}(\Omega))}\big)\|\bm{v}\|_{L^{2}(t_{1},+\infty;\bm{L}^{2}(\Omega))},$ (4.8)
@@
$\sup_{t\geq t_{1}(\varepsilon)}\frac{1}{2}\|\bm{v}(t)\|^{2}+\int_{t_{1}(\varepsilon)}^{+\infty}\!\int_{\Omega}2\nu(\varphi(s))|D\bm{v}(s)|^{2}\,\mathrm{d}x\mathrm{d}s\leq\frac{1}{2}\varepsilon^{2}+(1+|\chi|)\varepsilon^{2},$
@@
$\alpha\int_{0}^{+\infty}\!\!\int_{\Omega}|\overline{\varphi(s)}-c_{0}||\mu(s)|\,\mathrm{d}x\mathrm{d}s$ $\leq C\alpha\int_{0}^{+\infty}e^{-\alpha s}(1+\|\nabla\mu(s)\|)\,\mathrm{d}s$
$\leq C\alpha\left(1+\int_{0}^{+\infty}\|\nabla\mu(s)\|^{2}\,\mathrm{d}s\right)<+\infty,$
@@
By the Aubin–Lions compactness lemma, we infer from (4.4) and $\varphi\in H^{1}_{\mathrm{uloc}}([0,+\infty);H^{1}(\Omega))$ that $\varphi\in BUC([1,+\infty);W^{2-\epsilon,6}(\Omega))$ for any $\epsilon\in(0,1/2)$. In addition, due to the compactness embedding and uniqueness of the weak and strong limits, it holds $\varphi(t_{n})\to\varphi_{\infty}$ strongly in $W^{2-\epsilon,6}(\Omega)$ as $t_{n}\to+\infty$. Recalling the instantaneous regularity of $(\varphi,\mu,\sigma)$, for $t\geq 1$, we are allowed to test (2.2) by $\chi(\sigma-\chi\varphi)$ and (2.3) by $\sigma-\chi\varphi$, respectively. Adding the resultants together yields
@@
$\lim_{t_{n}\to+\infty}\|\sigma(t_{n})-[\chi\varphi_{\infty}+(\overline{\sigma_{0}}-\chi c_{0})]\|=0,$
and thus the limit $\sigma_{\infty}=\chi\varphi_{\infty}+(\overline{\sigma_{0}}-\chi c_{0})\in W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega)$.